# Lightweight Neural App Control

### v1

@@ Abstract
This paper introduces a novel mobile phone control architecture, termed “app agents”, for efficient interactions and controls across various Android apps. The proposed Lightweight Multi-modal App Control (LiMAC) takes as input a textual goal and a sequence of past mobile observations, such as screenshots and corresponding UI trees, to generate precise actions. To address the computational constraints inherent to smartphones, within LiMAC, we introduce a small Action Transformer (AcT) integrated with a fine-tuned vision-language model (VLM) for real-time decision-making and task execution. We evaluate LiMAC on two open-source mobile control datasets, demonstrating the superior performance of our small-form-factor approach against fine-tuned versions of open-source VLMs, such as Florence2 and Qwen2-VL. It also significantly outperforms prompt engineering baselines utilising closed-source foundation models like GPT-4o. More specifically, LiMAC increases the overall action accuracy by up to 19% compared to fine-tuned VLMs, and up to 42% compared to prompt-engineering baselines.
@@ 1 Introduction
footnotetext: *Equal contribution.footnotetext: Correspondence to: shaokun2@huawei.com
@@
Recent advancements have leveraged foundation models to develop app agents that understand natural language instructions and execute complex user commands within the smartphone’s interface (e.g.,  Rawles et al., 2024; Bai et al., 2024; Wang et al., 2024b; a). While foundation models enable sophisticated capabilities, relying on them for every action introduces significant drawbacks. Their substantial size and computational complexity make them resource-intensive and impractical for constant use on mobile devices. Alternatively, querying server-hosted foundation models, such as GPT-4o or Gemini, for each task can be prohibitively expensive due to the operational costs of running large models, making this approach impractical for everyday applications. For example, a state-of-the-art GPT-4o-based app agent (e.g.  Rawles et al., 2024) may require one to two minutes to run and cost approximately \$1.00 per task on average, based on tasks from the evaluated datasets.
@@ 2.2 Sequence modelling with transformers
Transformers (Vaswani et al., 2017) have demonstrated exceptional effectiveness in modelling and generating sequential data across a wide range of domains. They excel in various sequence modelling tasks, including those related to language, video processing, and decision-making (Chen et al., 2021). Regardless of the specific application, transformers begin by converting the input into a sequence of vectors. For text, this involves tokenising the input, with each token represented by an embedding vector. In the case of images, the input is typically divided into patches, where each patch is similarly represented by a vector, analogous to the tokenisation process in text. These embeddings, which map tokens or patches to vectors, can either be learned during the model’s training or sourced from pre-trained models (e.g.,  Devlin et al., 2018). However, one inherent limitation of the transformers is that treat the input as a set of unordered vectors, lacking any notion of temporal or spatial relationships. To address this, positional embeddings are added to the original token or patch embeddings to encode the position of each element in the sequence. By incorporating this positional information, the transformer is able to understand the relative ordering of the input.
The combined embeddings are fed through several multi-head self-attention layers, which are designed to capture dependencies and contextual relationships between different embeddings in the input sequence. These self-attention mechanisms allow the model to focus on relevant parts of the sequence when processing each embedding, enabling it to handle long-range dependencies more effectively. After passing through multiple layers, each consisting of self-attention and feed-forward components, the final activations from the transformer’s last hidden layer are passed through a linear (fully connected) layer. This layer is typically tasked with mapping the learned representations to the output space, whether for classification, prediction, or another specific task. The entire model is trained end-to-end, with backpropagation adjusting both the self-attention layers and the final linear layer to optimise performance on the desired task.
@@ 3 The lightweight multi-modal app control framework
Our methodology processes the user’s goal $g$ and the phone’s state at time $t$, utilising AcT, to determine the action type $a^{\text{type}}_{t}$. If the predicted action type is either input-text or open-app, then $g$, $o_{t}$, and $a^{\text{type}}_{t}$ are passed to a fine-tuned VLM, which is responsible for determining the specific action $a^{\text{spec}}_{t}$. For actions involving clicks, AcT handles the prediction directly but employs a different training objective that contrasts UI element embeddings to determine the most likely interaction target. Accordingly, this section is divided into three parts: predicting the action type, predicting specific actions for text input and app launching, and predicting clicks using our novel approach for interaction with UI elements. The full architecture of LiMAC is presented in Section 3.
@@ 3.1 Model inputs
UI Elements: The observed representation of each UI element $o_{t,i}$ at time $t$ is transformed into a vector $e^{\text{ui}}_{t,i}$ through several embedding functions. First, the text component is encoded using a sentence encoder (e.g., BERT) $e^{\text{txt}}_{t,i}=f_{\text{txt}}(o_{t,i}^{\text{txt}})$, and the image is encoded using a fine-tuned CLIP visual encoder (Radford et al., 2021) $e^{\text{img}}_{t,i}=f_{\text{img}}(o_{t,i}^{\text{img}})$. Additionally, any other attributes (e.g., clickable, editable, nested) are encoded into $e^{\text{attr}}_{t,i}=f_{\text{attr}}(o_{t,i}^{\text{attr}})$. The final embedding for each UI element is the concatenation of these vectors, $e^{\text{ui}}_{t,i}=[e^{\text{attr}}_{t,i};\,e^{\text{txt}}_{t,i};\,e^{\text{img}}_{t,i}]$. We fine-tune CLIP using the standard contrastive learning objective (Radford et al., 2021) using the screenshot of the observations and the related UI trees to allow adapting to app control datasets. We also add a positional encoding $p_{i}\in\mathbb{R}^{d}$ to represent the order or nesting of UI elements: $e^{\text{ui}}_{t,i}=e^{\text{ui}}_{t,i}+p_{i}$. This process is illustrated in Figure 1. To adapt the visual encoder $f_{\text{img}}$ to our task, we fine-tune it using our dataset by minimising the InfoNCE loss (Oord et al., 2018), aligning image and text representations of UI elements.
@@
Figure 2: The architecture of LiMAC. The history of observations-actions $\{o_{t},a_{t-1},o_{t-1}..\}$ and goal $g$ are processed to vector $x$ and passed to AcT. The image observation $o^{\text{img}}_{t}$ with the bounding boxes and the goal $g$ are passed as inputs to the VLM. The VLM is only called if an action that requires text completion is selected, based on the action type output of AcT. The action is finally selected based on the protocol described in Section 3.
@@ 3.3 Action type prediction
In our pipeline, the prediction of the next action begins with determining the action type. Predicting the action type $a_{t}^{\text{type}}$ can be framed as a classification problem, where we identify ten distinct action types in the datasets used in this work. These action types represent various possible interactions, such as click, open-app, scroll-down, input-text, or other essential commands. We implement the action type prediction with a specialised head. The action type head, denoted as $f_{\text{type}}$, transforms the final hidden state $h_{t}$ of the transformer (after the $e^{\text{end}}$ token) into a probability distribution over the possible action types, $p(a_{t}^{\text{type}}|h_{t})=f_{\text{type}}(h_{t})$. The learning objective for this task is to minimise the cross-entropy loss between the predicted and actual action types. Given a dataset $\mathcal{D}$, the cross-entropy loss for action type prediction is defined as:
@@ 3.4 Leveraging fine-tuned VLMs for text generation in action execution
As described in the previous section, our agent first predicts the action type. Among the ten action types, two specifically require textual specifications: i) the input-text action, where the specification is the text to be entered into a text box, and ii) the open-app action, where the specification is the name of the application to be opened. For these actions, we rely on fine-tuning a VLM using an app control dataset. The dataset provides action data in a dictionary-like format, such as: {"action-type":"open-app","app-name":"Chrome"}, with one key corresponding to the action type and another to the action specification. The VLM is trained to generate the correct sequence of tokens that corresponds to the successful completion of each action, optimising for the likelihood of generating the proper tokens based on the observation at each timestep.
During inference, after AcT predicts the action type, it guides the VLM by enforcing the model to start its response with the predicted action type. For instance, if AcT predicts input-text as the action type, the VLM is forced to begin its response with the following token pattern: {"action-type":"input-text","text":. The model then completes the specification, producing $a_{t}^{\text{spec}}$, which is the textual content needed for the action. The full action selection pipeline is presented in Figure 2.
@@ 4.1 Experimental setup
Datasets: Our experiments focus on two open mobile phone control datasets, AndroidControl (Li et al., 2024) and Android-in-the-Wild (AitW) (Rawles et al., 2023). Both datasets contain extensive human demonstrations of mobile phone navigation across a wide variety of tasks. In AndroidControl, every episode is defined by a specific goal, accompanied by a sequence of observations and actions. Each observation includes a screenshot from the phone and its corresponding UI tree. Conversely, observations in AitW lack the UI tree. As a result, it is necessary to extract the UI tree using an OCR system, which identifies all UI elements and provides a brief description of each. Further details on the goal format, observation space, and action space for each dataset can be found in Appendix A.
@@ 4.2 Evaluation pipeline
We evaluate on the test set of two datasets, using the same process for all models, with only the observation format and model calling differing. For each timestep, we call the model with the relevant observation format to generate an action. VLMs are trained to return actions in a specific format, while pre-trained models use a detailed prompt with the observation, as in Rawles et al. (2024).
One can calculate strict accuracy by directly comparing returned actions to the ground truth. However, in this work we relax this metric for a more practical assessment. In the relaxed version, a UI element is deemed correct if its bounding box is within the target element, as described by Li et al. (2024). For input-text actions, correctness is determined by a Jaccard index score of at least 0.5, reflecting the functional equivalence of similar inputs in search bars. We only report the relaxed accuracy metrics, found in Tables 1 and 2.
In Section 4.5 we also evaluate models’ action-type and click-target accuracy. Action-type accuracy reflects how well the model predicts the correct type for an action, regardless of the specifications such as the text content or target element. Click-target accuracy measures how accurately the model predicts the correct target for click actions when the action type is known. Computing the click-target accuracy requires rerunning a full evaluation over the dataset, where the output of the model is constrained to predict the click action and specify the target element. It should be noted that higher overall accuracy can still be achieved with lower action-type and/or click-target accuracy. This is because click-target accuracy is calculated separately, and because not all action types are equally advantageous for overall accuracy. Indeed, correctly predicting the type for actions that require additional specifications does not guarantee a positive overall outcome. Therefore, a method with slightly worse action-type accuracy, but prioritising actions without specifications, can achieve higher overall accuracy. For example, correctly predicting wait will directly translate to a successful timestep, while predicting input-text is still dependent on having the correct text.
@@ 4.4 Combining Different Modules
In the AndroidControl dataset, we observe that using M3A for predicting the target elements in click actions improves performance over using AcT alone. This demonstrates that GPT-4o is highly effective at identifying the correct target element when the prompt specifies that the action is click. This of courses comes at the cost of calling GPT-4o, which significantly increases the inference time. The highest overall accuracy is achieved when LiMAC is used to predict the action type, M3A is applied for target element prediction, and T3A is used for text generation In the AitW dataset, LiMAC combined with Florence for text generation yields the highest accuracy. This outcome is expected, as both M3A and T3A show significantly lower accuracy in this dataset (see Table 1).
@@ 4.5 Ablation Studies
Lastly, we present three ablation studies to further explore AcT design choices. A core feature of AcT is its ability to process each UI element as a distinct embedding within the transformer, created by concatenating the image, text, and attribute embeddings of the corresponding UI element. To assess the impact of the image and text modalities, as well as the CLIP fine-tuning on LiMAC’s performance, we compare it to three ablated versions: one that excludes the image component, another that omits the UI text in the embedding process, and one that uses the original CLIP for encoding the image embeddings instead of the fine-tuned version. The evaluation metrics for these comparisons in the AndroidControl dataset and using Florence2 for text completion are shown in Table 4.
The results demonstrate that removing image embeddings significantly reduces accuracy across all metrics, highlighting the crucial role of visual information in AcT. In contrast, omitting the text embeddings has only a slight effect on performance, suggesting that AcT can function effectively using only screenshots of observations without accessing the UI tree. Additionally, we observe that fine-tuning CLIP (see Section 3.1) is an important factor in improving the overall accuracy of LiMAC.
@@ 5 Related Work on App Control
Though graphical user interface (GUI) control mainly started with web-based datasets and foundation model agents (Shi et al., 2017; Liu et al., 2018; Yao et al., 2022a; Deng et al., 2023; Furuta et al., 2023; Gur et al., 2023; Zheng et al., 2024), there has recently been a significant focus on mobile phone control. This can be seen both by the rapid development of Android navigation datasets and environments (Rawles et al., 2023; 2024; Li et al., 2024), and of mobile control agents (Yang et al., 2023; Wang et al., 2024b; a; Wen et al., 2023; Hong et al., 2024; Rawles et al., 2024; Li et al., 2024; Bai et al., 2024). Though many agents are published with their own specific evaluation data, popular datasets such as Android-in-the-Wild (Rawles et al., 2023) or AndroidControl (Li et al., 2024) are often used as benchmarks.
The agents developed for this task can be divided into two clear input types: text-based, using UI accessibility tree or XML information to describe the screen, or image-based. Image-based agents require vision models, which are capable of directly processing image inputs, and are usually backed by VLMs. On the other hand, text-based agents are backed by classical LLMs. Image-based agents using VLMs also often take a combination of text and image as input to the model. Many mobile control agents propose intricate prompting methods backed by off-the-shelf, often proprietary, LLMs such as GPT-4 (Rawles et al., 2024; Yang et al., 2023; Wang et al., 2024b; a; Wen et al., 2023; Zheng et al., 2024). Though this requires little to no training, it can be both slow and expensive. Moreover, these models cannot be further tailored and trained for specific tasks. As such, another approach is to build mobile control agents around fine-tuned of foundation models on Android control datasets such as AitW or AndroidControl. Firstly, both AitW and AndroidControl present results for a fine-tuned LLM on their dataset, alongside the dataset itself. For example, Li et al. (2024) train various PaLM 2 (Anil et al., 2023) models on their dataset. However, these models are proprietary and supposedly quite large, with the base PaLM 2 model reported to have over 300B parameters. CogAgent (Hong et al., 2024) also performs fine-tuning on an 18B-large VLM. Bai et al. (2024) propose a different approach, called DigiRL, using RL to train their 1.3B VLM. This achieves strong performance but has its own limitations, notably the data gathering cost and simulation difficulty, leading to the model only being adept on a small subset of AitW.

### v2

@@ Abstract
This paper introduces a novel mobile phone control architecture, Lightweight Multi-modal App Control (LiMAC), for efficient interactions and control across various Android apps. LiMAC takes as input a textual goal and a sequence of past mobile observations, such as screenshots and corresponding UI trees, to generate precise actions. To address the computational constraints inherent to smartphones, we introduce a small Action Transformer (AcT) integrated with a fine-tuned vision-language model (VLM) for real-time decision-making and task execution. We evaluate LiMAC on two open-source mobile control datasets, demonstrating the superior performance of our small-form-factor approach against fine-tuned versions of open-source VLMs, such as Florence2 and Qwen2-VL. It also significantly outperforms prompt engineering baselines utilising closed-source foundation models like GPT-4o. More specifically, LiMAC increases the overall action accuracy by up to 19% compared to fine-tuned VLMs, and up to 42% compared to prompt-engineering baselines.
@@ 1 Introduction
footnotetext: * First authors with equal contribution.footnotetext: Corresponding author.
@@
Recent advancements have leveraged foundation models to develop app agents that understand natural language instructions and execute complex user commands within the smartphone’s interface (e.g.,  Rawles et al., 2024; Bai et al., 2024; Wang et al., 2024b; a). While foundation models enable sophisticated capabilities, relying on them for every action introduces significant drawbacks. Their substantial size and computational complexity make them resource-intensive and impractical for constant use on mobile devices. Alternatively, querying server-hosted foundation models, such as GPT-4o or Gemini, for each task can be prohibitively expensive due to the operational costs of running large models, making this approach impractical for everyday applications. For example, a state-of-the-art GPT-4o-based app agent (e.g.,  Rawles et al., 2024) may require one to two minutes to run and cost approximately \$1.00 per task on average, based on tasks from the evaluated datasets.
@@ 2.2 Sequence modelling with transformers
Transformers (Vaswani et al., 2017) have demonstrated exceptional effectiveness in modelling and generating sequential data across a wide range of domains. They excel in various sequence modelling tasks, including those related to language, video processing, and decision-making (Chen et al., 2021). Regardless of the specific application, transformers begin by converting the input into a sequence of vectors. For text, this involves tokenising the input, with each token represented by an embedding vector. In the case of images, the input is typically divided into patches, where each patch is similarly represented by a vector, analogous to the tokenisation process in text. These embeddings, which map tokens or patches to vectors, can either be learned during the model’s training or sourced from pre-trained models (e.g.,  Devlin et al., 2018). The embeddings are fed through several multi-head self-attention layers, which are designed to capture dependencies and contextual relationships between different embeddings in the input sequence. These self-attention mechanisms allow the model to focus on relevant parts of the sequence when processing each embedding, enabling it to handle long-range dependencies more effectively. After passing through multiple layers, each consisting of self-attention and feed-forward components, the final activations from the transformer’s last hidden layer are passed through a linear (fully connected) layer. This layer is typically tasked with mapping the learned representations to the output space, whether for classification, prediction, or another specific task. The entire model is trained end-to-end, with backpropagation adjusting both the self-attention layers and the final linear layer to optimise performance on the desired task.
@@ 3 The lightweight multi-modal app control framework
Our methodology processes the user’s goal $g$ and the phone’s state at time $t$, utilising AcT, to determine the action type $a^{\text{type}}_{t}$. If the predicted action type is either input-text or open-app, then $g$, $o_{t}$, and $a^{\text{type}}_{t}$ are passed to a fine-tuned VLM, which is responsible for determining the specific action $a^{\text{spec}}_{t}$. For actions involving clicks, AcT handles the prediction directly but employs a different training objective that contrasts UI element embeddings to determine the most likely interaction target. Accordingly, this section is divided into three parts: predicting the action type, predicting specific actions for text input and app launching, and predicting clicks using our novel approach for interaction with UI elements. The full architecture of LiMAC is presented below.
We refer to our method as lightweight because it uses fewer parameters on average during inference than baselines and, as we will show in Section 4.3, has faster inference speeds. The AcT module only has 520M parameters and the additional VLM component is called for less than 15% of actions in our datasets. LiMAC also selects actions more efficiently than a single VLM, as AcT does not require auto-regressive generation. While our approach has a higher memory footprint than solely using VLMs, due to loading both the AcT module and a VLM, its low parameter count remains within the capacity of modern devices (Li et al., 2024b; Laskaridis et al., 2024).
@@ 3.1 Model inputs
UI Elements: The observed representation of each UI element $o_{t,i}$ at time $t$ is transformed into a vector $e^{\text{ui}}_{t,i}$ through several embedding functions. First, the text component is encoded using a sentence encoder (e.g., BERT) $e^{\text{txt}}_{t,i}=f_{\text{txt}}(o_{t,i}^{\text{txt}})$, and the image is encoded using a fine-tuned CLIP visual encoder (Radford et al., 2021) $e^{\text{img}}_{t,i}=f_{\text{img}}(o_{t,i}^{\text{img}})$. Additionally, any other attributes (e.g., clickable, editable, nested) are encoded into $e^{\text{attr}}_{t,i}=f_{\text{attr}}(o_{t,i}^{\text{attr}})$. The final embedding for each UI element is the concatenation of these vectors, $e^{\text{ui}}_{t,i}=[e^{\text{attr}}_{t,i};\,e^{\text{txt}}_{t,i};\,e^{\text{img}}_{t,i}]$. We fine-tune CLIP using the standard contrastive learning objective (Radford et al., 2021) using the screenshot of the observations and the related UI trees to allow adapting to app control datasets. We also add a positional encoding $p_{i}\in\mathbb{R}^{d}$ to represent the order or nesting of UI elements: $e^{\text{ui}}_{t,i}=e^{\text{ui}}_{t,i}+p_{i}$. This process is illustrated in Figure 1. To adapt the visual encoder $f_{\text{img}}$ to our task, we fine-tune it using our dataset by minimising the InfoNCE loss (Oord et al., 2018), aligning image and text representations of UI elements. Similar methods of representing each UI element as an embedding for the transformer have been suggested by Li et al. (2020); Rawles et al. (2023), with the key distinction that our approach additionally fine-tunes the vision encoder to better adapt it for app control tasks.
@@
Figure 2: The architecture of LiMAC. The history of observations-actions $\{o_{t},a_{t-1},o_{t-1}..\}$ and goal $g$ are processed to vector $x$ and passed to AcT. The image observation $o^{\text{img}}_{t}$ with the bounding boxes and the goal $g$ are passed as inputs to the VLM. The VLM is only called if an action that requires text completion is selected, based on the action type output of AcT. The action is finally selected based on the protocol described in Sections 3.3, 3.4 and 3.5.
@@ 3.3 Action type prediction
In our pipeline, the prediction of the next action begins with determining the action type. Predicting the action type $a_{t}^{\text{type}}$ can be framed as a classification problem, where we identify a combined eleven distinct action types (see Appendix A), of which a subset are seen in individual datasets used in this work. These action types represent various possible interactions, such as click, open-app, scroll-down, input-text, or other essential commands. We implement the action type prediction with a specialised head. The action type head, denoted as $f_{\text{type}}$, transforms the final hidden state $h_{t}$ of the transformer (after the $e^{\text{end}}$ token) into a probability distribution over the possible action types, $p(a_{t}^{\text{type}}|h_{t})=f_{\text{type}}(h_{t})$. The learning objective for this task is to minimise the cross-entropy loss between the predicted and actual action types. Given a dataset $\mathcal{D}$, the cross-entropy loss for action type prediction is defined as:
@@ 3.4 Leveraging fine-tuned VLMs for text generation in action execution
As described in the previous section, our agent first predicts the action type. Among the eleven action types, two specifically require textual specifications: i) the input-text action, where the specification is the text to be entered into a text box, and ii) the open-app action, where the specification is the name of the application to be opened. For these actions, we rely on fine-tuning a VLM using an app control dataset. The dataset provides action data in a dictionary-like format, such as: {"action-type":"open-app","app-name":"Chrome"}, with one key corresponding to the action type and another to the action specification. The VLM is trained to generate the correct sequence of tokens that corresponds to the successful completion of each action, optimising for the likelihood of generating the proper tokens based on the observation at each timestep.
During inference, after predicting the action type, AcT guides the VLM to start its response with this action type. For instance, if AcT predicts input-text as the action type, the VLM is forced to begin its response with the token pattern: {"action-type":"input-text","text":. The model then completes the specification, producing $a_{t}^{\text{spec}}$, the textual content needed for the action. The full action selection pipeline is presented in Figure 2.
@@ 4.1 Experimental setup
Datasets: Our experiments focus on two open-source mobile phone control datasets, AndroidControl (Li et al., 2024a) and Android-in-the-Wild (AitW) (Rawles et al., 2023). Both contain extensive human demonstrations of mobile phone navigation across a wide variety of tasks. In AndroidControl, every episode is defined by a specific goal, accompanied by a sequence of observations and actions. Each observation includes a screenshot from the phone and its corresponding UI tree. Conversely, observations in AitW lack the UI tree. As a result, it is necessary to extract the UI tree using an OCR system that identifies all the UI elements and provides a brief description of each. More details on the goal format, observation space, and action space for each dataset can be found in Appendix A.
@@ 4.2 Evaluation pipeline
We evaluate on the test set of two datasets, using the same process for all models, with only the observation format and model calling differing. For each timestep, we call the model with the relevant observation format to generate an action. VLMs are trained to return actions in a specific format, while pre-trained models use a detailed prompt with the observation, as in Rawles et al. (2024). One can calculate strict accuracy by directly comparing returned actions to the ground truth. However, in this work we relax this metric for a more practical assessment, where a UI element is deemed correct if its bounding box is within the target element, as described by Li et al. (2024a). For input-text actions, correctness is determined by a Jaccard index score of at least 0.5, reflecting the functional equivalence of similar inputs in search bars. We report the relaxed accuracy metrics in Tables 1 and 2.
In Section 4.5 we also evaluate models’ action-type and click-target accuracy. Action-type accuracy reflects how well the model predicts the correct type for an action, regardless of the specifications such as the text content or target element. Click-target accuracy measures how accurately the model predicts the correct target for click actions when the action type is known. Computing the click-target accuracy requires rerunning a full evaluation over the dataset, where the output of the model is constrained to predict the click action and specify the target element. It should be noted that higher overall accuracy can still be achieved with lower action-type and/or click-target accuracy. This is because click-target accuracy is calculated separately, and because not all action types are equally advantageous for overall accuracy. Indeed, as defined in Section 2.1, an action is represented as $a_{t}=(a_{t}^{\text{type}},a_{t}^{\text{spec}})$, where both $a_{t}^{\text{type}}$ and $a_{t}^{\text{spec}}$ must be predicted correctly for a successful timestep. Actions which always have a null $a_{t}^{\text{spec}}$, like wait, are easier to predict correctly than those which have a complicated $a_{t}^{\text{spec}}$ that may be incorrectly predicted, like input-text.
@@ 4.4 Combining Different Modules
In the AndroidControl dataset, we observe that using M3A for predicting the target elements in click actions improves performance over using AcT alone. This demonstrates that GPT-4o is highly effective at identifying the correct target element when the prompt specifies that the action is click. This of courses comes at the cost of calling GPT-4o, which significantly increases the inference time. The highest overall accuracy is achieved when LiMAC is used to predict the action type, M3A is applied for target element prediction, and T3A is used for text generation. In the AitW dataset, LiMAC combined with Florence for text generation yields the highest accuracy. This outcome is expected, as both M3A and T3A show significantly lower accuracy in this dataset (see Table 1).
@@ 4.5 Ablation Studies
Lastly, we present three ablation studies to further explore AcT design choices. A core feature of AcT is its ability to process each UI element as a distinct embedding within the transformer, created by concatenating the image, text, and attribute embeddings of the corresponding UI element. To assess the impact of the image and text modalities, as well as the CLIP fine-tuning on LiMAC’s performance, we compare it to three ablated versions: one that excludes the image component, another that omits the UI text in the embedding process, and one that uses the original CLIP for encoding the image embeddings instead of the fine-tuned version. The evaluation metrics for these comparisons in the AndroidControl dataset and using Florence2 for text completion are shown in Table 4. The results demonstrate that removing image embeddings significantly reduces accuracy across all metrics, highlighting the crucial role of visual information in AcT. In contrast, omitting the text embeddings has only a slight effect on performance, suggesting that AcT can function effectively using only screenshots of observations without accessing the UI tree. Additionally, we observe that fine-tuning CLIP (see Section 3.1) is an important factor in improving the overall accuracy of LiMAC.
@@ 5 Related Work on App Control
Though graphical user interface (GUI) control mainly started with web-based datasets and foundation model agents (Shi et al., 2017; Liu et al., 2018; Yao et al., 2022a; Deng et al., 2023; Furuta et al., 2023; Gur et al., 2023; Zheng et al., 2024), there has recently been a significant focus on mobile phone control. This can be seen both by the rapid development of Android navigation datasets, environments, and benchmarks (Rawles et al., 2023; 2024; Li et al., 2024a; Chen et al., 2024), and of mobile control agents (Yang et al., 2023; Wang et al., 2024b; a; Wen et al., 2023; Hong et al., 2024; Rawles et al., 2024; Li et al., 2024a; Bai et al., 2024; Wang et al., 2024c). Though many agents are published with their own specific evaluation data, popular datasets such as Android-in-the-Wild (Rawles et al., 2023) or AndroidControl (Li et al., 2024a) are often used as benchmarks. Agents developed for this task can be divided into two clear input types: text-based, using UI accessibility tree or XML information to describe the screen, or image-based. Image-based agents require vision models, which are capable of directly processing image inputs, and are usually backed by VLMs. On the other hand, text-based agents are backed by classical LLMs. Image-based agents also often take a combination of text and image as input to the model. Many mobile control agents propose intricate prompting methods backed by off-the-shelf, often proprietary, LLMs such as GPT-4 (Rawles et al., 2024; Yang et al., 2023; Wang et al., 2024b; a; Wen et al., 2023; Zheng et al., 2024). Although this requires little to no training, it can be both slow and expensive. Moreover, these models cannot be further tailored and trained for specific tasks. As such, another approach is to build mobile control agents around fine-tuned of foundation models on Android control datasets such as AitW or AndroidControl. Firstly, both AitW and AndroidControl present results for a fine-tuned LLM on their dataset, alongside the dataset itself. For example, Li et al. (2024a) train various PaLM 2 (Anil et al., 2023) models on their dataset. However, these models are proprietary and supposedly quite large, with the base PaLM 2 model reported to have over 300B parameters. CogAgent (Hong et al., 2024) also performs fine-tuning on an 18B-large VLM. Bai et al. (2024) propose a different approach, called DigiRL, using RL to train their 1.3B VLM. This achieves strong performance but has limitations such as gathering cost and simulation difficulty, leading to the model only being adept on a small subset of AitW.